if you combine 370.0 mL of water at 25.00 ∘C and 140.0 mL of water at 95.00 ∘C, what is the final temperature of the mixture? Use 1.00 g/mL as the density of water.
1
Expert's answer
2019-10-17T07:18:08-0400

The problem can be solved by treating the problem as a weighted average":"

So,


"T_f=\frac{T_{1}V_1+T_2V_2}{V_1+V_2}=\frac{(370\times 25)+(140\times 95)}{370+95}=48.49\ \degree\ C"

So,Final temprature =


"T_f=48.49\ \degree C"